### v2

@@ 4 Proof of Theorem 4
3             if $v$ does not satisfy any of (c1), (c2) and (c3) in Lemma 7 then return ‘NO’ if $v$ satisfies (c1) or (c3) in Lemma 7 then
4                   let $A$ be the connected component of $G[L_{i-1}\setminus Q]$ s.t. $N(v)\cap A=\{u\}$ $\phi\leftarrow\phi\cup\{(v\lor u),(\neg v\lor\neg u)\}$ $Q\leftarrow Q\cup\{v,u\}$ foreach $x\in N(v)\setminus Q$ do  $\phi\leftarrow\phi\cup\{(v\lor\neg x),(\neg v\lor x)\}$ foreach $x\in N(u)\setminus Q$ do  $\phi\leftarrow\phi\cup\{(u\lor\neg x),(\neg u\lor x)\}$
@@
7                   let $u_{1},u_{2}$ be the two neighbors of $v$ in $L_{i-1}\setminus Q$, and let $w\in L_{i-2}\setminus Q$ be such that $v,u_{1},u_{2},w$ induce a $4$-cycle $\phi\leftarrow\phi\cup\{(v\lor w),(\neg v\lor\neg w),(u_{1}\lor u_{2}),(\neg u_{1}\lor\neg u_{2})\}$ $Q\leftarrow Q\cup\{v,u_{1},u_{2},w\}$ foreach $x\in N(v)\setminus Q$ do  $\phi\leftarrow\phi\cup\{(v\lor\neg x),(\neg v\lor x)\}$ foreach $x\in N(w)\setminus Q$ do  $\phi\leftarrow\phi\cup\{(w\lor\neg x),(\neg w\lor x)\}$ foreach $x\in N(u_{1})\setminus Q$ do  $\phi\leftarrow\phi\cup\{(u_{1}\lor\neg x),(\neg u_{1}\lor x)\}$ foreach $x\in N(u_{2})\setminus Q$ do  $\phi\leftarrow\phi\cup\{(u_{2}\lor\neg x),(\neg u_{2}\lor x)\}$
@@
if $\phi$ is satisfiable then return $(X,Y)$ with $X$ the true and $Y$ the false vertices else return ‘NO’
Algorithm 1 Recognizing $4$-chordal graphs having pmc
@@
In Fig. 6, consider the graph on the left. Letting $r=v_{2}$, we have $L_{0}=\{v_{2}\}$, $L_{1}=\{v_{0},v_{1},v_{3}\}$ and $L_{2}=\{v_{4},v_{5}\}$. Depending on the choice of $v\in L_{3}\setminus Q$ in line 1 of the algorithm, we have two cases:
$v_{1}$$v_{0}$$v_{4}$$v_{5}$$v_{2}$$v_{3}$$v_{2}$$v_{0}$$v_{1}$$v_{3}$$v_{5}$$v_{4}$
@@
  • Suppose the algorithm first picks $v=v_{4}$. Then the formula $\phi$ consists of the following clauses: $(v_{4}\lor v_{3})$, ($(\neg v_{4}\lor\neg v_{3})$ (line 1), $(v_{4}\lor\neg v_{5})$, ($(\neg v_{4}\lor v_{5})$, $(v_{3}\lor\neg v_{5})$, ($(\neg v_{3}\lor v_{5})$, $(v_{3}\lor\neg v_{2})$, ($(\neg v_{3}\lor v_{2})$ (lines 1, 1), $(v_{5}\lor v_{1})$, ($(\neg v_{5}\lor\neg v_{1})$ (line 1), $(v_{1}\lor\neg v_{0})$, ($(\neg v_{1}\lor v_{0})$, $(v_{1}\lor\neg v_{2})$, ($(\neg v_{1}\lor v_{2})$ (lines 1, 1), and $(v_{0}\lor v_{2})$, ($(\neg v_{0}\lor\neg v_{2})$ (line 1). In this case, the output is ‘NO’ as $\phi$ is unsatisfiable (line 1).
  • Suppose the algorithm first picks $v=v_{5}$. Then the formula $\phi$ consists of the following clauses: $(v_{5}\lor v_{2})$, ($(\neg v_{5}\lor\neg v_{2})$, $(v_{1}\lor v_{3})$, ($(\neg v_{1}\lor\neg v_{3})$ (line 1), $(v_{5}\lor\neg v_{4})$, ($(\neg v_{5}\lor v_{4})$, $(v_{2}\lor\neg v_{0})$, ($(\neg v_{2}\lor v_{0})$, $(v_{1}\lor\neg v_{0})$, ($(\neg v_{1}\lor v_{0})$, $(v_{3}\lor\neg v_{4})$, ($(\neg v_{3}\lor v_{4})$ (lines 11). In this case, the output is ‘NO’ as in the next round $v=v_{4}\in L_{2}\setminus Q$ does not satisfy any of the conditions (c1), (c2) and (c3) (line 1).
@@
  • First let $r=v_{0}$, implying $L_{0}=\{v_{0}\}$, $L_{1}=\{v_{1},v_{2}\}$, $L_{2}=\{v_{3},v_{4}\}$ and $L_{3}=\{v_{5}\}$. Then the formula $\phi$ consists of the following clauses: $(v_{5}\lor v_{1})$, ($(\neg v_{5}\lor\neg v_{1})$, $(v_{3}\lor v_{4})$, ($(\neg v_{3}\lor\neg v_{4})$ (line 1), $(v_{1}\lor\neg v_{0})$, ($(\neg v_{1}\lor v_{0})$, $(v_{3}\lor\neg v_{2})$, ($(\neg v_{3}\lor v_{2})$ (lines 11), and $(v_{2}\lor v_{0})$, ($(\neg v_{2}\lor\neg v_{0})$ (line 1). In this case, $\phi$ has a satisfying assignment, say $v_{0}=v_{1}=v_{4}=\text{True}$ and $v_{2}=v_{3}=v_{5}=\text{False}$, and the output is the perfect matching cut $(X,Y)$ with $X=\{v_{0},v_{1},v_{4}\}$ and $Y=\{v_{2},v_{3},v_{5}\}$ (line 1).
  • Next let $r=v_{1}$, implying $L_{0}=\{v_{1}\}$, $L_{1}=\{v_{0},v_{3},v_{4}\}$ and $L_{2}=\{v_{2},v_{5}\}$. By symmetry, we may assume that the algorithm first picks $v=v_{2}\in L_{2}\setminus Q$ at line 1. Then the formula $\phi$ consists of the following clauses: $(v_{2}\lor v_{1})$, ($(\neg v_{2}\lor\neg v_{1})$, $(v_{0}\lor v_{3})$, ($(\neg v_{0}\lor\neg v_{3})$ (line 1), $(v_{1}\lor\neg v_{4})$, ($(\neg v_{1}\lor v_{4})$, $(v_{3}\lor\neg v_{5})$, ($(\neg v_{4}\lor v_{5})$ (lines 11), and $(v_{5}\lor v_{4})$, ($(\neg v_{5}\lor\neg v_{4})$ (line 1). Again, $\phi$ has a satisfying assignment, say $v_{0}=v_{1}=v_{4}=\text{True}$ and $v_{2}=v_{3}=v_{5}=\text{False}$, and the output is the perfect matching cut $(X,Y)$ with $X=\{v_{0},v_{1},v_{4}\}$ and $Y=\{v_{2},v_{3},v_{5}\}$ (line 1).
@@
Observe that $\phi$ has $O(|E|)$ clauses. Since checking the conditions (c1), (c2) and (c3) in Lemma 7, as well as deciding if a 2sat formula is satisfiable and if so computing a satisfying assignment AspvallPT79 , can be done in polynomial time, the running time of Algorithm 1 clearly is polynomial.
@@
We prove this fact by induction on $i$. First consider the case $i=h$. In this case, any $v\in L_{h}\setminus Q$ is a leaf (as $L_{h+1}=\emptyset$). Hence, by Lemma 7, every $v\in L_{h}\setminus Q$ must satisfy (c1), (c2) or (c3), otherwise $G$ does have no perfect matching cuts and the output NO at line 1 is correct. Thus, suppose that $G$ has a perfect matching cut.
@@
Suppose first that $\phi$ admits a truth assignment. Then the partition $(X,Y)$ of $V(G)$ with $X$ the set of all ‘true’ vertices and $Y$ the set of all ‘false’ vertices is a perfect matching cut of $G$: the clauses constructed at lines 1 and 1 ensure that $X$ and $Y$ are non-empty, and any vertex in $X$ has a neighbor in $Y$ and vice versa. Moreover, the clauses constructed at lines 1 and 1, and at lines 11 ensure that if $x\in X$ (respectively $x\in Y$) has a neighbor $p(x)\in Y$ (respectively $p(x)\in X$) then all neighbors $p(x)$ of $x$ belong to the same part with $x$ and all neighbors $x$ of $p(x)$ belong to the same part with $p(x)$. That is, $x$ has exactly on neighbor $p(x)$ in the other part.
Conversely, suppose that $G$ has a perfect matching cut $(X,Y)$. Then set variable $v$ to True if the corresponding vertex $v$ is in $X$, and False if it is in $Y$. With Lemma 7 we argue that $\phi$ is satisfied by this truth assignment: the two clauses at line 1 are satisfied because in this case $u$ and $v$ must belong to different parts, and also the clauses at lines 1 and 1 because the other neighbors $x$ of $v$ and $u$ unless their private neighbors must belong to the same part with $v$, respectively, with $u$. Similar, the four clauses at line 1 are satisfied because in this case $v$ and $w$, as well as $u_{1}$ and $u_{2}$ must belong to different parts, and also the clauses at lines 11 because the other neighbors $x$ of $v,w,u_{1}$ and $u_{2}$ unless their private neighbors must belong to the same part with $v,w,u_{1}$ and $u_{2}$, respectively.
@@ 5 Conclusions
We have shown that all three problems matching cut, perfect matching cut and disconnected perfect matching are $\mathsf{NP}$-complete in $P_{14}$-free $8$-chordal graphs. The hardness result for perfect matching cut solves an open problem posed in FeghaliLPR22 ; LuckePR22isaac ; LuckePR22arxiv ; LuckePR23J . For matching cut and disconnected perfect matching, the hardness result improves the previously results in LuckePR23J in $P_{15}$-free graphs, respectively, in $P_{19}$-free graphs, to $P_{14}$-free graphs. An obvious question is whether one of these problems remains $\mathsf{NP}$-complete in $P_{t}$-free graphs for some $t<14$.
We have proved that disconnected perfect matching and perfect matching cut are solvable in polynomial time when restricted to $4$-chordal graphs. Thus, another obvious question is whether one of these problems remains $\mathsf{NP}$-complete in $k$-chordal graphs for $5\leq k\leq 7$.

### v3

@@ 6 $Q\leftarrow\emptyset$; $\phi\leftarrow\emptyset$
10    if $v$ does not satisfy any of (c1), (c2) and (c3) in Lemma 7 then return ‘NO’
11    if $v$ satisfies (c1) or (c3) in Lemma 7 then
12         let $A$ be the connected component of $G[L_{i-1}\setminus Q]$ s.t. $N(v)\cap A=\{u\}$
13       $\phi\leftarrow\phi\cup\{(v\lor u),(\neg v\lor\neg u)\}$
14       $Q\leftarrow Q\cup\{v,u\}$
15       foreach $x\in N(v)\setminus Q$ do $\phi\leftarrow\phi\cup\{(v\lor\neg x),(\neg v\lor x)\}$
16       foreach $x\in N(u)\setminus Q$ do $\phi\leftarrow\phi\cup\{(u\lor\neg x),(\neg u\lor x)\}$
17
@@ 14       $Q\leftarrow Q\cup\{v,u\}$
20         let $u_{1},u_{2}$ be the two neighbors of $v$ in $L_{i-1}\setminus Q$, and let $w\in L_{i-2}\setminus Q$ be such that $v,u_{1},u_{2},w$ induce a 44-cycle
21       $\phi\leftarrow\phi\cup\{(v\lor w),(\neg v\lor\neg w),(u_{1}\lor u_{2}),(\neg u_{1}\lor\neg u_{2})\}$
22       $Q\leftarrow Q\cup\{v,u_{1},u_{2},w\}$
23       foreach $x\in N(v)\setminus Q$ do $\phi\leftarrow\phi\cup\{(v\lor\neg x),(\neg v\lor x)\}$
24       foreach $x\in N(w)\setminus Q$ do $\phi\leftarrow\phi\cup\{(w\lor\neg x),(\neg w\lor x)\}$
25       foreach $x\in N(u_{1})\setminus Q$ do $\phi\leftarrow\phi\cup\{(u_{1}\lor\neg x),(\neg u_{1}\lor x)\}$
26       foreach $x\in N(u_{2})\setminus Q$ do $\phi\leftarrow\phi\cup\{(u_{2}\lor\neg x),(\neg u_{2}\lor x)\}$
27
@@ 22       $Q\leftarrow Q\cup\{v,u_{1},u_{2},w\}$
33if $\phi$ is satisfiable then return $(X,Y)$ with $X$ the true and $Y$ the false vertices
34 else return ‘NO’
35
Algorithm 1 Recognizing 44-chordal graphs having pmc
@@
In Fig. 6, consider the graph on the left. Letting $r=v_{2}$, we have $L_{0}=\{v_{2}\}$, $L_{1}=\{v_{0},v_{1},v_{3}\}$ and $L_{2}=\{v_{4},v_{5}\}$. Depending on the choice of $v\in L_{2}\setminus Q$ in line 1 of the algorithm, we have two cases:
$v_{1}$$v_{0}$$v_{4}$$v_{5}$$v_{2}$$v_{3}$
$v_{2}$$v_{0}$$v_{1}$$v_{3}$$v_{5}$$v_{4}$
@@
  • Suppose the algorithm first picks $v=v_{4}$. Then the formula $\phi$ consists of the following clauses: $(v_{4}\lor v_{3})$, $(\neg v_{4}\lor\neg v_{3})$ (line 1), $(v_{4}\lor\neg v_{5})$, $(\neg v_{4}\lor v_{5})$, $(v_{3}\lor\neg v_{5})$, $(\neg v_{3}\lor v_{5})$, $(v_{3}\lor\neg v_{2})$, $(\neg v_{3}\lor v_{2})$ (lines 1, 1), $(v_{5}\lor v_{1})$, $(\neg v_{5}\lor\neg v_{1})$ (line 1), $(v_{1}\lor\neg v_{0})$, $(\neg v_{1}\lor v_{0})$, $(v_{1}\lor\neg v_{2})$, $(\neg v_{1}\lor v_{2})$ (lines 1, 1), and $(v_{0}\lor v_{2})$, $(\neg v_{0}\lor\neg v_{2})$ (line 1). In this case, the output is ‘NO’ as $\phi$ is unsatisfiable (line 1).
  • Suppose the algorithm first picks $v=v_{5}$. Then the formula $\phi$ consists of the following clauses: $(v_{5}\lor v_{2})$, $(\neg v_{5}\lor\neg v_{2})$, $(v_{1}\lor v_{3})$, $(\neg v_{1}\lor\neg v_{3})$ (line 1), $(v_{5}\lor\neg v_{4})$, $(\neg v_{5}\lor v_{4})$, $(v_{2}\lor\neg v_{0})$, $(\neg v_{2}\lor v_{0})$, $(v_{1}\lor\neg v_{0})$, $(\neg v_{1}\lor v_{0})$, $(v_{3}\lor\neg v_{4})$, $(\neg v_{3}\lor v_{4})$ (lines 11). In this case, the output is ‘NO’ as in the next round $v=v_{4}\in L_{2}\setminus Q$ does not satisfy any of the conditions (c1), (c2) and (c3) (line 1).
@@
  • First let $r=v_{0}$, implying $L_{0}=\{v_{0}\}$, $L_{1}=\{v_{1},v_{2}\}$, $L_{2}=\{v_{3},v_{4}\}$ and $L_{3}=\{v_{5}\}$. Then the formula $\phi$ consists of the following clauses: $(v_{5}\lor v_{1})$, $(\neg v_{5}\lor\neg v_{1})$, $(v_{3}\lor v_{4})$, $(\neg v_{3}\lor\neg v_{4})$ (line 1), $(v_{1}\lor\neg v_{0})$, $(\neg v_{1}\lor v_{0})$, $(v_{3}\lor\neg v_{2})$, $(\neg v_{3}\lor v_{2})$ (lines 11), and $(v_{2}\lor v_{0})$, $(\neg v_{2}\lor\neg v_{0})$ (line 1). In this case, $\phi$ has a satisfying assignment, say $v_{0}=v_{1}=v_{4}=\text{True}$ and $v_{2}=v_{3}=v_{5}=\text{False}$, and the output is the perfect matching cut $(X,Y)$ with $X=\{v_{0},v_{1},v_{4}\}$ and $Y=\{v_{2},v_{3},v_{5}\}$ (line 1).
  • Next let $r=v_{1}$, implying $L_{0}=\{v_{1}\}$, $L_{1}=\{v_{0},v_{3},v_{4}\}$ and $L_{2}=\{v_{2},v_{5}\}$. By symmetry, we may assume that the algorithm first picks $v=v_{2}\in L_{2}\setminus Q$ at line 1. Then the formula $\phi$ consists of the following clauses: $(v_{2}\lor v_{1})$, $(\neg v_{2}\lor\neg v_{1})$, $(v_{0}\lor v_{3})$, $(\neg v_{0}\lor\neg v_{3})$ (line 1), $(v_{1}\lor\neg v_{4})$, $(\neg v_{1}\lor v_{4})$, $(v_{3}\lor\neg v_{5})$, $(\neg v_{3}\lor v_{5})$ (lines 11), and $(v_{5}\lor v_{4})$, $(\neg v_{5}\lor\neg v_{4})$ (line 1). Again, $\phi$ has a satisfying assignment, say $v_{0}=v_{1}=v_{4}=\text{True}$ and $v_{2}=v_{3}=v_{5}=\text{False}$, and the output is the perfect matching cut $(X,Y)$ with $X=\{v_{0},v_{1},v_{4}\}$ and $Y=\{v_{2},v_{3},v_{5}\}$ (line 1).
@@
Observe that $\phi$ has $O(|E|)$ clauses. Since checking the conditions (c1), (c2) and (c3) in Lemma 7, as well as deciding if a 2sat formula is satisfiable and if so computing a satisfying assignment [1], can be done in polynomial time, the running time of Algorithm 1 clearly is polynomial.
@@
We prove this fact by induction on $i$. First consider the case $i=h$. In this case, any $v\in L_{h}\setminus Q$ is a leaf (as $L_{h+1}=\emptyset$). Hence, by Lemma 7, every $v\in L_{h}\setminus Q$ must satisfy (c1), (c2) or (c3), otherwise $G$ has no perfect matching cuts and the output NO at line 1 is correct. Thus, suppose that $G$ has a perfect matching cut.
@@
Suppose first that $\phi$ admits a truth assignment. Then the partition $(X,Y)$ of $V(G)$ with $X$ the set of all ‘true’ vertices and $Y$ the set of all ‘false’ vertices is a perfect matching cut of $G$: the clauses constructed at lines 1 and 1 ensure that $X$ and $Y$ are non-empty, and any vertex in $X$ has a neighbor in $Y$ and vice versa. Moreover, the clauses constructed at lines 1 and 1, and at lines 11 ensure that if $x\in X$ (respectively $x\in Y$) has a neighbor $p(x)\in Y$ (respectively $p(x)\in X$) then all neighbors $\not=p(x)$ of $x$ belong to the same part with $x$ and all neighbors $\not=x$ of $p(x)$ belong to the same part with $p(x)$. That is, $x$ has exactly one neighbor $p(x)$ in the other part.
Conversely, suppose that $G$ has a perfect matching cut $(X,Y)$. Then set variable $v$ to True if the corresponding vertex $v$ is in $X$, and False if it is in $Y$. With Lemma 7, we argue that $\phi$ is satisfied by this truth assignment: the two clauses at line 1 are satisfied because in this case $u$ is the private neighbor of $v$ and $v$ is the private neighbor of $u$, hence $u$ and $v$ must belong to different parts. Also, the clauses at lines 1 and 1 are satisfied because the other neighbors $x$ of $v$, respectively, of $u$ must belong to the same part as $v$, respectively, as $u$. Similar, the four clauses at line 1 are satisfied because in this case $v$ and $w$, as well as $u_{1}$ and $u_{2}$ must belong to different parts. Finally, the clauses at lines 11 are satisfied because the other neighbors $x$ of $v,w,u_{1}$ and $u_{2}$ must belong to the same part as $v,w,u_{1}$ and $u_{2}$, respectively.
@@ 5 Conclusions
We have shown that all three problems matching cut, perfect matching cut and disconnected perfect matching are $\mathsf{NP}$-complete in $P_{14}$-free $8$-chordal graphs. The hardness result for perfect matching cut in $P_{14}$-free graphs solves an open problem posed in [22]. For matching cut and disconnected perfect matching, the hardness result improves the previous results in [22] in $P_{15}$-free graphs, respectively, in $P_{19}$-free graphs, to $P_{14}$-free graphs. An obvious question is whether one of these problems remains $\mathsf{NP}$-complete in $P_{t}$-free graphs for some $t<14$.
The hardness result for perfect matching cut in $8$-chordal graphs partly solves an open problem posed in [19, 23]. We have proved that disconnected perfect matching and perfect matching cut are solvable in polynomial time when restricted to $4$-chordal graphs. Thus, another obvious question is whether one of these problems remains $\mathsf{NP}$-complete in $k$-chordal graphs for $5\leq k\leq 7$.
Acknowledgments
We would like to thank one of the reviewers for her/his extremely careful reading and many detailed and useful comments.